C&A’s potato chip filling process has a lower specification limit of 9.5 oz. and an upper specification limit of 10.5 oz. The standard deviation is 0.3 oz. and the mean is 10 oz. What is the probability that a bag will weigh more than 10.5 oz.?

95.2%

90.4%

9.56%

4.78%
Answer #1

1. D. 4.78


In order to find the probability, we will need the following formula:

Z = MEAN - USL / STDEV * SQRT(NO. OF UNITS)

SINCE WE HAVE TO FIND THE PROBABILITY OF 1 BAG GOING ABOVE THE USL, WE WILL CALCULATE THE OTHER VALUES:

Z = 10 - 10.5 / 0.3 = -1.666666667

USING THE NORMAL TABLE, WE ASSOCIATE THE VALUE TO REPRESENT 0.0477 OR 4.78%(ROUNDED TO 2 DECIMAL PLACES)